OpenStudy (anonymous):

add and write in simplest form: 4y/y+2+8/y+2

OpenStudy (anonymous):

4

OpenStudy (hoblos):

is it like this? 4y 8 --------- + --------------- y+2 y+2

OpenStudy (anonymous):

\[\frac{4y}{y+2}+\frac{8}{y+2}\] \[\frac{4y+8}{y+2}=\frac{4(y+2)}{y+2}=4\]

OpenStudy (anonymous):

yes it is:)